OpenStudy (anonymous):

Which ratio is proportional to 10/24>?

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{10}{24}= \frac{10 \color{red} {\times x } }{24\color{red} {\times x } } }\)

OpenStudy (solomonzelman):

whether x is a fraction or a whole number. (no mater what x is)

OpenStudy (anonymous):

> that is not suppose to be there

OpenStudy (solomonzelman):

I figured :)

OpenStudy (solomonzelman):

you see what I mean. you can divide top and bottom by any number (as long it's divided by same thing on top as on bottom) and get a ratio that's equivalent.

OpenStudy (solomonzelman):

like 5/12 (for instance)

OpenStudy (anonymous):

so 5/8

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

\(\huge\color{green}{ \frac{10}{24} }\) \(\huge\color{green}{ =\frac{10\color{red} {\div 2 } }{24\color{red} {\div2 } } =\frac{?}{?} }\) you tell me

OpenStudy (anonymous):

5/12

OpenStudy (anonymous):

there is no 5/12

OpenStudy (solomonzelman):

what other options do you have ?

OpenStudy (anonymous):

A-35/84 B-5/6 C-20/36 D5/8

OpenStudy (solomonzelman):

you know that B and D don't work, because the only proportion with a nominator of 5, is with denominator of 12. right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

10/24 does C work? checking... 20/36 to get 20 you multiply ten TIMES TWO but if you multiply 24 times 2, you DO NOT get 36

OpenStudy (anonymous):

i see

OpenStudy (solomonzelman):

and to get a, multiply top and bottom times 7/2